Make New Entry,
Make Followup Entry
User name dhamilton
Log entry time 00:03:28 on November21,2008
Entry number 250114
keyword=Swing Shift Summary
Swing Shift Summary
David Hamilton (SL)
John LeRose (TO)
Xiaoyan Deng (3rd)
16:00 - Shift starts with beam tuning and bringing up HRSl magnets following down-time today.
17:45 - BPMs, harp scans, raster are all fine. Begin production data-taking on 3He target.
18:25 - Stopped run 3770 early because deadtime was at 100%. Restarted RunControl. Everything fine with 3771.
22:10 - DAQ crash. Had to reboot ROCs 5, 9 and 10.
VESDA sensor readings have been steadily incresing from around 7pm. MCC and Dave Kausch know of the problem - it's the dump dehumidifier.
Run 3770 17:45 0.3M - stopped early due to DAQ problem (100% dead)
Run 3771 18:30 1.0M
Run 3772 19:40 1.0M
Run 3773 20:45 1.0M
Run 3774 21:45 0.3M - stopped early for DAQ restart (ROC5).
Run 3775 22:20 0.8M - stopped early to restart HAPPEX run.
Run 3776 23:12 .... still running.